# LaZagna: An Open-Source Framework for Flexible 3D FPGA Architectural Exploration

### v1

@@ Abstract
While 3D IC technology has been extensively explored for ASICs, their application to FPGAs remains limited. Existing studies on 3D FPGAs are often constrained to fixed prototypes, narrow architectural templates, and simulation-only evaluations. In this work, we present LaZagna, the first open-source framework for automated, end-to-end 3D FPGA architecture generation and evaluation. LaZagna supports high-level architectural specification, synthesizable RTL generation, and bitstream production, enabling comprehensive validation of 3D FPGA designs beyond simulation. It significantly broadens the design space compared to prior work by introducing customizable vertical interconnect patterns, novel 3D switch block designs, and support for heterogeneous logic layers. The framework also incorporates practical design constraints such as inter-layer via density and vertical interconnect delay. We demonstrate the capabilities of LaZagna by generating synthesizable RTL that can be taken through full physical design flows for fabric generation, along with functionally correct bitstreams. Furthermore, we conduct five case studies that explore various architectural parameters and evaluate their impact on wirelength, critical path delay, and routing runtime. These studies showcase the framework’s scalability, flexibility, and effectiveness in guiding future 3D FPGA architectural and packaging decisions. LaZagna is fully open-source and available on GitHub [^1] https://anonymous.4open.science/r/LaZagna-784F.
@@ I Introduction
II Previous Works and Limitations
@@
TABLE I: LaZagna’s capabilities in modeling a wide range of prior 3D FPGA architectures.
Related works fall into two main categories: (1) architectural studies on 3D FPGAs, and (2) tools for FPGA fabric and bitstream generation.
3D FPGA Architecture Studies. Various strategies have been proposed for vertical integration and layer organization in 3D FPGAs. These architectures can be categorized into the following classes:
@@
Non-logic heterogeneous designs have also demonstrated great performance potential. Amagasaki et al. [8] also explored a type 2 architecture that separates logic and routing across layers using 3D switch blocks, achieving a 5.3% CPD reduction over a 2D baseline. Turkyilmaz et al. [4] showed a 14% CPD improvement by separating configuration memory from logic. Lin et al. [5] isolated configuration memory onto a dedicated layer, resulting in an estimated 1.7$\times$ CPD improvement. Waqar et al. [6] further advanced this idea using back-end-of-line integration to place configuration memory and SBs above CLBs, reporting CPD improvements ranging from 11% to 30%.
@@ III Tool Workflow
In LaZagna, these patterns are user-defined. Both the input and output patterns are expressed as a sequence of four integers corresponding to the four planar sides of the switch block in counter-clockwise order. Each integer indicates the starting track index on that side to be used in vertical connectivity. Consider a switch block with channel width $W$ and an output pattern defined as $[i_{0},i_{1},i_{2},i_{3}]$. The $k$-th output cross-layer track will be driven by the planar tracks indexed as $(i_{0}+k)\bmod W$ on the left side, $(i_{1}+k)\bmod W$ on the bottom, $(i_{2}+k)\bmod W$ on the right, and $(i_{3}+k)\bmod W$ on the top, where $k$ ranges from 00 to $W-1$. The input pattern operates analogously, except that the specified planar track indices are the targets driven by each vertical input track.
@@ III-A6 All Standard 2D FPGA Parameters
  • 3D SB Signals. Each 3D SB must also correctly identify inter-layer signals from other 3D SBs. To send an inter-layer signal, the SB multiplexes input routing tracks based on the SB pattern, marking the output as a inter-layer signal. To receive an inter-layer signal, the SB inspects other SBs with the same location but on different layers. If an inter-layer output is detected, its added to the SB and the SB de-multiplexes and routes it to its output tracks based on the output pattern.
@@ III-D Benchmark Logic Synthesis, Placement, and Routing
With the 3D-RRG generated, the benchmark circuit undergoes logic synthesis with Yosys (performed independently of 3D-RRG), followed by benchmark PnR onto the 3D fabric using VTR. Since the 3D-RRG reflects the architecture of the target FPGA, VTR can apply its PnR algorithms directly and produce valid routing results. These results include performance metrics such as CPD, WL, and resource utilization. We note that a key limitation of VTR’s 3D placement algorithm is that logic blocks are assigned to specific layers during initial placement and are not allowed to move across layers thereafter. This constraint may lead to suboptimal placements. Moreover, VTR does not currently support heterogeneous 3D architectures or CB-based vertical connectivity for architectures with more than two layers. Addressing these limitations remains an avenue for future toolchain enhancements.
@@ III-F Tool Output
As discussed earlier, LaZagna produces three primary outputs: the RTL for the 3D FPGA fabric, benchmark PnR results, and the corresponding bitstream. The RTL enables downstream analysis and physical design, allowing designers detailed power, performance, area, and thermal (PPAT) evaluations, and potentially GDSII layouts for 3D FPGAs.
@@ IV-B Tool Demonstration
To verify both the functionality of the generated 3D FPGA RTL and the validity of the bitstream generated, the s298 MCNC benchmark [15] was implemented onto the homogeneous and non-logic heterogeneous fabrics. Bitstream generation is not currently possible for the logic heterogeneous architectures, as explained in Section III-C. The bitstream was generated using the flow described in Section III. The fabrics were then functionally verified by programming them with the bitstream in RTL simulation and comparing their outputs to the expected results. This process confirmed the correctness of both the generated RTL and bitstream.
@@ IV-C1 Impact of Vertical Connection Types
WL and CPD. Fig. 6 and Table IV present the CPD and WL results of our study. All 3D architectures significantly outperform the 2D baseline, with reductions in CPD ranging from 17.9% to 21.9% and 19.0% to 25.3% for WL. The 3D Hybrid architectures exhibit the best performance, with the 100% 3D SB configuration achieving the largest WL decrease and the 66% 3D SB configuration showing the largest CPD reduction. There is a clear correlation between the number of vertical connections and performance improvement.
Takeaways. (1) 3D FPGAs show promising improvements in CPD and WL compared to 2D FPGAs. (2) Increasing the number of vertical interconnects generally leads to better performance in 3D FPGAs. (3) A balanced architecture that combines different vertical connection types tends to yield the most effective trade-offs.
@@
Vertical Connection Type # of Vert. Connections Vert. Conns. Per Grid WL Geomean (reduc.) CPD Geomean (reduc.)
3D CB 3,038,192 316.3 0.754 (24.6%) 0.799 (20.1%)
3D CB-O 431,984 45.0 0.810 (19.0%) 0.810 (19.0%)
3D SB 100% 750,680 78.2 0.752 (24.8%) 0.821 (17.9%)
3D SB 66% 495,598 51.6 0.753 (24.7%) 0.819 (18.1%)
3D SB 33% 247,832 25.8 0.755 (24.5%) 0.809 (19.1%)
3D Hybrid 100% 3,788,872 394.5 0.747 (25.3%) 0.782 (21.8%)
3D Hybrid 66% 3,533,790 367.9 0.752 (24.8%) 0.781 (21.9%)
3D Hybrid 33% 3,286,064 342.2 0.751 (24.9%) 0.788 (21.2%)
3D Hybrid-O 100% 1,182,664 123.1 0.769 (23.1%) 0.794 (20.6%)
3D Hybrid-O 66% 927,582 96.6 0.768 (23.2%) 0.790 (21.0%)
3D Hybrid-O 33% 679,816 70.8 0.770 (23.0%) 0.785 (21.5%)
@@ IV-C2 Impact of 3D SB Placement
This case study investigates the impact of 3D SB placement on CPD, WL, and routing time. Building on the promising results of the 3D SB connection type from Case Study 1, we explore if the spatial distribution of 3D SBs has a significant effect on performance. We use 3D SB subset connections with 50% 3D SBs and evaluate six placement strategies, illustrated in Fig.2(b). The corresponding results are presented in Fig.7.
Among the strategies, the perimeter placement emerges as a clear outlier, exhibiting significantly worse CPD and WL than the other approaches, and even under-performing the 2D baseline in terms of CPD. Additionally, it incurs an exponentially higher routing time. This is likely due to the VTR placement engine being unaware of vertical connectivity, potentially placing logic blocks that require vertical links far from the perimeter-placed 3D SBs, thereby increasing routing congestion and runtime. These observations highlight the need for 3D-aware placement algorithms.
Excluding the perimeter case, the remaining strategies exhibit comparable CPD and WL. However, routing time varies. Strategies with more distributed 3D SBs (Repeated Interval and Random) achieve the fastest routing times, while those with concentrated 3D SBs (Core and Perimeter) perform substantially worse. This suggests that a more uniform distribution of 3D SBs across the fabric can lead to more efficient routing.
@@ IV-C3 Impact of 3D SB Connection Patterns
Results shown in Fig. 8 indicate that WL is largely unaffected by the SB pattern. CPD shows a small variation, with improvements ranging from 16.9% to 18.9%. The revolving input and subset patterns yielded the best CPD results. While further investigation is needed to understand the impact of different patterns, these initial results suggest that the placement of 3D SBs has a greater influence than the connection pattern.
Takeaways. (1) 3D SB connection patterns appear to have a minimal effect on WL and a relatively small impact on CPD. (2) The location of 3D SBs appears to be a more significant factor.
@@ IV-C4 Impact of Vertical Connection Delay
This case study examines the effect of vertical connection delay on 3D FPGA performance. Different 3D integration technologies exhibit varying vertical connection densities and speeds. To explore this impact, we varied the vertical delay ratio from 0.5$\times$ to 10$\times$ the delay of a length-4 horizontal routing segment. As a reference point, [10] used a vertical connection delay of 137ps was used (0.739$\times$ the horizontal segment delay), representing hybrid bonding with a 5µm pitch. In our study, a 0.5$\times$ vertical delay corresponds to a vertical delay of 93ps, while a 10$\times$ ratio corresponds to 1.86ns. This wide range encompasses various technologies, from faster hybrid bonding (93ps or less depending on pitch) to slower TSVs (e.g., 1.5ns as reported in [19]).
Fig. 9 presents the results. CPD exhibits a near-linear relationship with the vertical delay ratio. WL increases slightly for 3D CB and 3D CB-O architectures but remains almost constant for 3D SB architectures. However, a key limitation in VTR’s placement engine may influence these results. VTR assigns blocks to layers pseudo-randomly during initial placement, with a preference for layers containing connected blocks. Crucially, blocks cannot be moved between layers after this step.
This fixed layer assignment potentially limits performance gains, especially considering the interplay between vertical delay and layer crossings. As shown in Fig. 10, the number of layer crossings remains constant for 3D CB-O architectures and decreases slightly for 3D SB and 3D CB architectures as the vertical delay ratio increases. This is a direct consequence of VTR’s placement limitations. Ideally, with a low vertical delay, designs should maximize layer crossings to exploit vertical connectivity. Conversely, a high vertical delay should minimize crossings. VTR’s fixed layer assignment prevents such dynamic optimization.
Takeaways. (1) CPD increases approximately linearly with vertical connection delay. However, VTR’s placement limitations may influence this relationship. (2) 3D FPGAs demonstrate performance comparable to their 2D counterparts when the vertical connection delay is approximately five times that of a horizontal connection. This suggests that even with moderately high vertical delays, 3D architectures can remain competitive.
@@ IV-C5 Impact of Benchmark Placement Algorithm
The result reveals that 3D SB architectures generally achieve better CPD across all benchmarks. However, it also shows that the variability in CPD across different seeds is comparable between 2D and 3D architectures. For instance, the conv_layer_hls benchmark exhibits a wide range of CPD results depending on the seed, in contrast to the robot_rl benchmark, which demonstrates a significantly tighter distribution. This suggests that the introduction of 3D interconnect does not inherently increase the variability of placement and routing results. However, the benchmark-specific variability warrants further investigation.
@@ V Conclusion and Future Work
In this work, we introduced LaZagna, an automated tool that streamlines the design and exploration of 3D FPGA architectures. By generating custom 3D FPGA fabrics based on user-defined parameters, LaZagna facilitates efficient exploration of the vast 3D FPGA design space. And allows the exploration of 3D FPGAs beyond simulation by producing synthesizable RTL that can be physically designed. Our results demonstrate that 3D FPGAs can achieve improvements of up to 21.9% for critical path delay and up to 25.3% for total wire length compared to 2D FPGAs.

### v3

@@ Abstract
While 3D IC technology has been extensively explored for ASICs, their application to FPGAs remains limited. Existing studies on 3D FPGAs are often constrained to fixed prototypes, narrow architectural templates, and simulation-only evaluations. In this work, we present LaZagna, the first open-source framework for automated, end-to-end 3D FPGA architecture generation and evaluation. LaZagna supports high-level architectural specification, synthesizable RTL generation, and bitstream production, enabling comprehensive validation of 3D FPGA designs beyond simulation. It significantly broadens the design space compared to prior work by introducing customizable vertical interconnect patterns, novel 3D switch block designs, and support for heterogeneous logic layers. The framework also incorporates practical design constraints such as inter-layer via density and vertical interconnect delay. We demonstrate the capabilities of LaZagna by generating synthesizable RTL that can be taken through full physical design flows for fabric generation, along with functionally correct bitstreams. Furthermore, we conduct five case studies that explore various architectural parameters and evaluate their impact on wirelength, critical path delay, and routing runtime. These studies showcase the framework’s scalability, flexibility, and effectiveness in guiding future 3D FPGA architectural and packaging decisions. LaZagna is fully open-source and available on GitHub [^1] https://github.com/sharc-lab/LaZagna
@@ I Introduction
II Previous Work and Limitations
@@
TABLE I: LaZagna’s 3D FPGA modeling capabilities compared to previous works.
Related work falls into two main categories: (1) architectural studies on 3D FPGAs, and (2) tools for FPGA fabric and bitstream generation.
3D FPGA Architecture Studies. Various strategies have been proposed for vertical integration and layer organization in 3D FPGAs. We categorize these architectures into the following classes:
@@
[…]
@@ III Tool Workflow
In LaZagna, these patterns are user-defined. Both the input and output patterns are expressed as a sequence of four integers corresponding to the four planar sides of the switch block in counter-clockwise order. Each integer indicates the starting track index on that side to be used in vertical connectivity. Consider a switch block with channel width $W$ and an output pattern defined as $i_{0},i_{1},i_{2},i_{3}]$. The $k$-th output cross-layer track will be driven by the planar tracks indexed as $(i_{0}+k)\bmod W$ on the left side, $(i_{1}+k)\bmod W$ on the bottom, $(i_{2}+k)\bmod W$ on the right, and $(i_{3}+k)\bmod W$ on the top, where $k$ ranges from 0 to W1W-1. The input pattern operates analogously, except that the specified planar track indices are the targets driven by each vertical input track.
@@ III-A6 All Standard 2D FPGA Parameters
  • 3D SB Signals. Each 3D SB must also correctly identify inter-layer signals from other 3D SBs. To send an inter-layer signal, the SB multiplexes input routing tracks based on the SB pattern, marking the output as an inter-layer signal. To receive an inter-layer signal, the SB inspects other SBs with the same location but on different layers; if an inter-layer output is detected, the SB de-multiplexes it based on its output pattern and routes it to the appropriate output tracks.
@@ III-D Benchmark Logic Synthesis, Placement, and Routing
With the 3D-RRG generated, the benchmark circuit undergoes logic synthesis with Yosys (performed independently of 3D-RRG), followed by benchmark PnR onto the 3D fabric using VTR. Since the 3D-RRG reflects the architecture of the target FPGA, VTR can apply its PnR algorithms directly and produce valid routing results. These results include performance metrics such as CPD, WL, and resource utilization. We note that a key limitation of VTR’s 3D placement algorithm is that logic blocks are assigned to specific layers during initial placement and are rarely allowed to move across layers thereafter. This constraint may lead to suboptimal placements. Moreover, VTR does not currently support heterogeneous 3D architectures or CB-based vertical connectivity for architectures with more than two layers. Addressing these limitations remains an avenue for future toolchain enhancements.
@@ III-F Tool Output
LaZagna produces three primary outputs: the RTL for the 3D FPGA fabric, benchmark PnR results, and the corresponding bitstream. The RTL enables downstream analysis and physical design, allowing designers detailed power, performance, area, and thermal (PPAT) evaluations, and potentially GDSII layouts for 3D FPGAs.
@@ IV-B Tool Demonstration
To verify both the functionality of the generated 3D FPGA RTL and the validity of the bitstream generated, the s298 MCNC benchmark [15] was implemented onto the homogeneous and non-logic heterogeneous fabrics. Bitstream generation is not currently possible for the logic heterogeneous architectures, as explained in Section III-D. The bitstream was generated using the flow described in Section III. The fabrics were then functionally verified by programming them with the bitstream in RTL simulation and comparing their outputs to the expected results. This process confirmed the correctness of both the generated RTL and bitstream.
@@ IV-C1 Impact of Vertical Connection Types
WL and CPD. Figure 6 and Table IV summarize the WL and CPD results. All 3D architectures outperform the 2D baseline, achieving CPD reductions of 3.1%–6.7% and WL reductions of 4.2%–22.9%. Among them, the 3D Hybrid architectures deliver the most balanced improvements in both metrics, with the 3D CB architecture achieving the largest WL reduction and the 100% 3D SB configuration yielding the greatest CPD reduction. While WL improvement shows a clear positive correlation with the number of vertical connections, CPD does not follow the same trend: the 3D SB architectures achieve the best CPD reductions despite having relatively few vertical connections per grid.
Takeaways. (1) 3D FPGAs show promising improvements in CPD and WL compared to 2D FPGAs. (2) While more vertical interconnects improve WL performance, the relationship with CPD is less straightforward, with diminishing returns at higher connection densities. (3) A balanced architecture that combines different vertical connection types tends to yield the most effective trade-offs.
@@
Vertical Connection Type # of Vert. Connections Vert. Conns. Per Grid WL Geomean (% reduc.) CPD Geomean (% reduc.)
3D CB 3,038,192 316.3 0.771 (22.9%) 0.965 (3.5%)
3D CB-O 431,984 45.0 0.958 (4.2%) 0.969 (3.1%)
3D SB 100% 750,680 78.2 0.884 (11.6%) 0.933 (6.7%)
3D SB 66% 495,598 51.6 0.882 (11.8%) 0.950 (5.0%)
3D SB 33% 247,832 25.8 0.890 (11.0%) 0.943 (5.7%)
3D Hybrid 100% 3,788,872 394.5 0.827 (17.3%) 0.957 (4.3%)
3D Hybrid 66% 3,533,790 367.9 0.818 (18.2%) 0.951 (4.9%)
3D Hybrid 33% 3,286,064 342.2 0.826 (17.4%) 0.960 (4.0%)
3D Hybrid-O 100% 1,182,664 123.1 0.906 (9.4%) 0.956 (4.4%)
3D Hybrid-O 66% 927,582 96.6 0.916 (8.4%) 0.955 (4.5%)
3D Hybrid-O 33% 679,816 70.8 0.917 (8.3%) 0.960 (4.0%)
@@ IV-C2 Impact of 3D SB Placement
This case study investigates the impact of 3D SB placement on CPD, WL, and routing time. Building on the promising results of the 3D SB connection type from Case Study 1, we explore if the spatial distribution of 3D SBs has a significant effect on performance. We use 3D SB subset connections with 50% 3D SBs and evaluate six placement strategies, illustrated in Fig.2(b). The corresponding results are presented in Fig.7. Among the evaluated strategies, the Perimeter placement stands out as a clear outlier, yielding substantially worse CPD and WL than all other approaches and even performing significantly below the 2D baseline. It also incurs an exponentially higher routing time. This degradation is likely due to the VTR placement engine’s lack of awareness of vertical connectivity, which may result in logic blocks requiring vertical links being placed far from the perimeter-located 3D SBs, thereby increasing routing congestion and prolonging runtime. These findings underscore the importance of developing 3D-aware placement algorithms.
Excluding the Perimeter case, the remaining strategies exhibit comparable CPD, WL, and routing times, with the Core placement showing a slightly higher runtime than the others. Strategies with more distributed 3D SBs, such as Repeated Interval, Rows, Columns, and Random, achieve the fastest routing times, while those with concentrated 3D SBs, such as Core and Perimeter, perform worse. This suggests that a more uniform distribution of 3D SBs across the fabric can enable more efficient routing.
@@ IV-C3 Impact of 3D SB Connection Patterns
Results shown in Fig. 8 indicate that WL and CPD are largely unaffected by the SB pattern. CPD and WL show small variation, with CPD improvements ranging from 4.3% to 6.7%, and WL improvements ranging from 10.4% to 12.1%. The symettric offset and subset patterns yielded the best CPD results. While further investigation is needed to understand the impact of different patterns, these initial results suggest that the placement of 3D SBs has a greater influence than the connection pattern.
Takeaways. (1) 3D SB connection patterns appear to have a minimal effect on WL and CPD. (2) The location of 3D SBs appears to be a more significant factor.
@@ IV-C4 Impact of Vertical Connection Delay
[…]
Fig. 9 presents the results. CPD exhibits a near-linear relationship with the vertical delay ratio. WL increases slightly for all 3D architectures. However, a key limitation in VTR’s placement engine may influence these results. During initial placement, VTR assigns blocks to layers pseudo-randomly, preferring layers containing connected blocks. Subsequently, the algorithm limits inter-layer movement, with only a $\sim$10% probability of layer reassignment per block during the placement flow.
This limited layer reassignment potentially limits performance gains, especially considering the interplay between vertical delay and layer crossings. As shown in Fig. 10, the number of layer crossings remains constant for 3D CB-O architectures and decreases slightly for 3D SB and 3D CB architectures as the vertical delay ratio increases. This is a direct consequence of VTR’s placement limitations. Ideally, with a low vertical delay, designs should maximize layer crossings to exploit vertical connectivity. Conversely, a high vertical delay should minimize crossings. VTR’s limited layer reassignment prevents such dynamic optimization.
Takeaways. (1) CPD increases approximately linearly with vertical connection delay. However, VTR’s placement limitations may influence this relationship. (2) 3D FPGAs demonstrate performance comparable to their 2D counterparts when the vertical connection delay is approximately two times that of a horizontal connection. This suggests that even with moderately high vertical delays, 3D architectures can remain competitive.
@@ IV-C5 Impact of Benchmark Placement Algorithm
The results indicate that 3D SB architectures consistently achieve better CPD across all benchmarks. However, the spread of CPD values across different placement seeds is similar for both 2D and 3D architectures. For example, the attention_layer benchmark shows a wide variation in CPD depending on the seed, whereas the conv_layer benchmark exhibits a much tighter distribution. This suggests that introducing 3D interconnects does not inherently increase the variation in placement and routing outcomes. Nonetheless, the benchmark-specific differences in CPD variability merit further investigation.
@@ V Conclusion and Future Work
In this work, we introduced LaZagna, an automated tool that streamlines the design and exploration of 3D FPGA architectures. By generating custom 3D FPGA fabrics based on user-defined parameters, LaZagna facilitates efficient exploration of the vast 3D FPGA design space. And allows the exploration of 3D FPGAs beyond simulation by producing synthesizable RTL that can be physically designed. Our results demonstrate that 3D FPGAs can achieve improvements of up to 6.7% for critical path delay and up to 22.9% for total wire length compared to 2D FPGAs.